A spring extends 3 inches when loaded vertically with 4 lbs. This weight was pulled further down by 3 inches an was released from rest. Determine the frequency of vibration in rad per second


1
Expert's answer
2021-04-21T07:04:14-0400

4lbs = 1.81 kgs

3inches = 0.0762 m

"k= \frac{17.8Newtons}{ 0.0762m}=233.59N/m"

= 237.5328N/m

"\omega_n= \sqrt{\frac{k}{m}}= \sqrt{\frac{233.59}{1.81}}=11.35 rad/s"